# Feasibility of Federated Learning from Client Databases with Different Brain Diseases and MRI Modalities

### v1

@@ 3.1 Datasets & Prepocessing
We use 7 annotated brain MRI databases, each with different MRI modalities and segmentation tasks collected from different sites: ATLAS v2.0 [13] for stroke lesion, BRATS 2016 [2] for tumor, MSSEG [3] for multiple sclerosis, TBI from our site for traumatic brain injury, WMH [10] for White Matter Hyperintensity, ISLES 2015’s SISS subtask [14] for stroke lesion, and Tumor2 for tumour segmentation from our institution. The data is split into training and validation data. Tab. 1 gives an overview of the data sizes and modalities. We resample the images to 1x1x1 mm, skull-strip and apply z-score intensity normalization, and combine all lesion types across databases into a single ’lesion’ label.
Table 1: MRI databases overview: ISLES and Tumor2 are used for evaluation only, except for single center training, where a 70-30 train-validation split is used.
Database Modalities Disease/Task #Samples #Val. Samples
T1 T1c FLAIR T2 PD SWI
BRATS Tumor 484 40
MSSEG Multiple Sclerosis 53 16
ATLAS Stroke Lesion 654 195
TBI Traumatic Brain Injury 281 125
WMH White Matter Hyperintensity 60 18
ISLES Stroke Lesion 28 28
Tumor2 Tumor 57 57
@@ 3.2 Experimental Setup
Setup & Methods: We use the Residual U-Net [29] as network architecture and Adam optimizer. We linearly scale the learning rate from 0 to $10^{-3}$ over 50 communication rounds, keep it constant until communication round 150, and then linearly decay it. We set $\tau=58$ (equivalent to 1 epoch for the largest database, ATLAS), $E=300$ (or 300 epochs for centralised training), and use a batch size of 8. The Dice score on the validation set is reported after the last epoch. For our federated experiments, both the standard FedAvg algorithm [15] and, separately, FedBN [12] (to keep client-specific BN layers) are used. As discussed in Section 2.2, in addition to BN, our experiments investigate the impact of alternative feature normalization methods that do not rely on client-specific normalization layers: Group Norm (GN) [25], Instance Norm (IN) [23] and the NF [32] technique.
@@ 3.3 Results & Discussion
Unified Model - Improves Performance Across Databases: Results in Tab. 2 show that, when keeping client-specific BN parameters, federated joint training across diverse MRI databases with different brain diseases and different input modalities improves the model’s performance on each database, either surpassing or matching single center training. This highlights the benefits of transferring knowledge between different brain diseases, especially for databases with limited data (MSSEG, TBI, and WMH). Training with modality drop and client-specific BN layers leads to the highest overall results and improves over FedAvg using averaged BN parameters, demonstrating the benefits of client specific BN layers for source clients. Furthermore, this joint learning framework produces a single model instead of training 5 individual models, demonstrating its practicality.
Table 2: Separate models are trained and evaluated on each database (single center), with only BN results shown as they yield the best results. In the federated setting the model is trained across ATLAS, BRATS, MSSEG, TBI, and WMH with various feature normalization methods. Furthermore, results of modality drop are reported.
Setting Mod.Drop Norm ATLAS BRATS MSSEG TBI WMH Average
Single Database BN 52.8 91.9 68.3 56.1 71.5 68.1
Federated (client spec. BN params) BN 50.9 91.8 70.2 55.8 74.8 68.7
Single Database BN 52.8 91.7 66.7 53.9 69.1 66.8
Federated (avg. BN params) BN 43.1 89.4 63.2 51.4 67.3 62.9
Federated IN 49.0 90.7 66.6 52.9 71.0 66.0
Federated GN 48.1 91.0 67.3 53.7 70.4 66.1
Federated NF 51.2 91.1 64.8 54.4 72.3 66.8
Federated (client spec. BN params) BN 54.5 91.8 69.1 56.2 73.7 69.1
Modality Drop Improves Robustness Against Missing Modalities: Our initial step towards developing a model capable of segmenting pathologies with modality combinations different from those seen during training involves robustness against missing modalities. Results in Tab. 3 show the need for modality drop to improve the model’s robustness against missing modalities. When using modality drop during training, the Dice score decreases by only 3.9 when evaluating with missing modalities compared to all modalities, versus a 21.3 decrease without modality drop. Without modality drop, the model relies on having all modalities for a specific disease present during inference. Therefore, the following experiments focus on models trained with modality drop.
Table 3: Results from jointly training a model using client-specific BN layers on ATLAS, BRATS, MSSEG, TBI, and WMH, both with and without modality drop and evaluated on the same databases by randomly excluding modalities during testing to mimic missing modalities, as well as with all modalities present during testing. Training with modality drop is important so that at test-time the model can work with different combinations of modalities. The average Dice score across all databases is reported.
Training Method Evaluation (Dice) $\Delta$ Dice
Modality Drop No Modality Drop
No Modality Drop 47.4 68.7 21.3
Modality Drop 65.2 69.1 03.9
Zero-Shot Generalization to Unseen Modality Combinations: Tab. 4 presents generalization results on unseen databases with modality combinations not seen during training. It is remarkable that despite not seeing any of the new client’s data, the best federated model’s average Dice score reaches only 4.1% below the upper bound of single center training on the target database and comes close to centralized joint training (for which only results with IN are shown as it achieved best performance). Traditionally, models are trained on databases specific to a disease and evaluated on unseen clients with the same disease, as we trained with ATLAS (for stroke lesion) and evaluated on ISLES, and trained with BRATS (for brain tumors) and evaluated on Tumor2, setting this as baseline. Both are only trained and evaluated on T1 since it is the only overlapping modality. Federated joint training achieves a performance gain of 49.3 on ISLES and 8.1 on Tumor2, reinforcing the benefits of joint training across multiple MRI databases compared to pathology specific models.
Additionally, modality drop proves essential to dissociate specific modality combinations with certain pathologies, as all jointly trained models with modality drop show improved average performance on segmenting pathologies with modality combinations not seen during training. For Tumor2, there is extensive improvement (both in centralised and federated settings), whereas for ISLES, the federated models performed unexpectedly slightly worse, likely due to gradient stochasticity in federated optimization.
Comparing client-specific BN layers to non-client-specific normalization approaches, the results indicate that the NF technique outperforms all other techniques. Using the average of client-specific BN layers and estimating the running statistics on the validation data gives the second-best results. However, this estimation is less practical as it requires unlabeled data from the unseen client.
@@
Table 4: Generalization results on unseen clients ISLES and Tumor2, each with modality combinations not seen during training (reported in Dice score): Separate models were trained on ISLES and Tumor2 and evaluated on the same database (Upper Bound). Baselines involve training on stroke lesion database ATLAS, evaluating on ISLES, and training on tumor database BRATS, evaluating on Tumor2. These results are compared to FL, using different normalization techniques, across 5 training databases (ATLAS, BRATS, MSSEG, TBI, and WMH) and evaluated on ISLES and Tumor2. Centralised training on all 5 databases together is also presented.
@@
Setting Mod. Drop Norm. Needs Targ. Data ISLES Tumor2 Average
BN N/A 57.9 75.3 66.6
IN N/A 55.7 76.7 66.2
Upper Bound: Trained on Target only NF N/A 60.6 72.6 66.6
Trained on ATLAS (T1) IN N/A 5.3 -
Trained on BRATS (T1) IN N/A - 63.7 34.5
Centralised IN No 50.7 58.1 54.4
Centralised IN No 55.5 72.2 63.9
Federated (avg. BN params) BN No 53.8 0.0 26.9
Federated (avg. BN params) BN No 43.1 68.0 55.6
Federated IN No 48.3 59.1 53.7
Federated IN No 44.2 72.7 58.5
Federated NF No 54.6 31.1 42.9
Federated NF No 52.8 72.1 62.5
Federated (client spec. BN params) BN Yes 50.2 36.8 43.5
Federated (client spec. BN params) BN Yes 49.9 70.1 60.0
@@
Investigating the Effects of Normalization Layers on Source Clients: Tab. 2 shows the effects of feature normalization on the source databases. The findings indicate that keeping client-specific BN layers yields better performance compared to other normalization techniques on the source clients, but pose difficulties when dealing with unseen clients (as shown in Table 4). The NF method outperforms other non client-specific feature normalization layers in both source client and generalization results. Consequently, a trade-off must be made when designing a model for training with this type of heterogeneous data: either improve the source clients’ performance or prioritizing the development of a model that generalizes to unseen clients.
@@ 4 Conclusion
This study shows the feasibility, practicality and effectiveness of using FL to train a single model across decentralised brain MRI databases, each with different input modalities and segmentation tasks. Our results show that federated joint training on multiple databases results in a single model that can segment all diseases and sets of modalities seen in training clients as well as client-specific models trained for a specific set of modalities and disease. Furthermore, it shows promising generalization to unseen clients with modality combinations not seen during training, and highlight the importance of the feature normalization method’s choice to address data heterogeneity. We believe this study highlights the benefits of leveraging heterogeneous brain MRI databases for federated training and opens up future research directions in this area.

### v2

@@ 4.1 Datasets & Prepocessing
We use 7 annotated brain lesion MRI databases, each with different MRI modalities and segmentation tasks, collected from different sites: ATLAS v2.0 [23] for stroke lesion, BRATS [3] via the Medical Segmentation Decathlon [2] for tumor, MSSEG [10] for multiple sclerosis, TBI from our site for traumatic brain injury, WMH [19] for White Matter Hyperintensity, ISLES 2015’s SISS subtask [26] for stroke lesion, and Tumor2 for tumour segmentation from our institution. The dataset is divided into training/validation sets as follows: BRATS (444/40), MSSEG (37/16), ATLAS (459/195), TBI (156/125), WMH (42/18), ISLES (20/8), and Tumor2 (40/17). For the zero-shot generalization experiments (Sec. 4.3, Tab. 4), all samples are used to evaluate the model. For the Oracle results (Tab. 4), where the model is trained on the target database and not all samples can be used for evaluation, we apply 3-fold cross-validation, ensuring each sample is used for evaluation once. Tab. 3 gives an overview of the available modalities. We resample the images to 1x1x1 mm, skull-strip, apply z-score intensity normalization, and combine all lesion types across databases into a single ‘lesion’ label.
Table 3: MRI databases overview. Shown are the sets of modalities in each database and the corresponding segmentation tasks.
Database Modalities Disease/Task
T1 T1c FLAIR T2 PD SWI DWI
BRATS Tumor
MSSEG Multiple Sclerosis
ATLAS Stroke Lesion
TBI Traumatic Brain Injury
WMH White Matter Hyperintensity
ISLES Stroke Lesion
Tumor2 Tumor
@@ 4.2 Experimental Setup
All experiments use the Residual U-Net [45] as backbone, Adam optimizer, and $\alpha\!=\!0.8$ for our weighted loss function. We linearly scale the learning rate from 0 to $10^{-3}$ over 50 communication rounds, keep it constant until communication round 150, and then linearly decay it. We set $\tau\!=\!58$ (equivalent to 1 epoch for the largest database, ATLAS), $E\!=\!300$ (or 300 epochs for centralised training), and batch size of 8. The Dice score on the validation set is reported after the last epoch. For Federated Learning, we study the effect of feature normalisation methods, assessing BN when all its parameters are aggregated as per the standard FedAvg algorithm [27], and, separately, using FedBN [21] to keep client-specific BN parameters (cf. Sec. 3.2). We also assess methods that do not rely on client-specific normalization layers: GN [40], IN [38], and NF [48].
@@ 4.3 Results & Discussion
Unified Model Improves Performance Across Databases
@@
Experimental results in Tab. 2 prove that it is feasible to use Federated Learning to train a single model jointly over client databases with diverse diseases and sets of modalities and reach performance levels of database-specific models. When learning client-specific BN parameters, FedUniBrain matches or even surpasses single center training, highlighting the benefits of knowledge transfer between brain diseases. Training with modality drop and client-specific BN layers leads to the highest overall results and performs better than using averaged BN parameters, demonstrating the benefits of client specific BN layers for source clients.  Visual examples of training a single model per client compared to FedUniBrain are shown in Fig. 3(a). Besides its practicality over disease-specific models, the unified model may have improved generalisation abilities on unseen databases, which are assessed in follow-up experiments.
@@
Table 4: Assessing generalization to unseen databases, ISLES and Tumor2, each with modality combinations not seen during training (Dice score % reported): Separate models were trained on ISLES and Tumor2 and evaluated on the same database as Oracle baselines. We train task-specific baselines on stroke database ATLAS, then evaluate on ISLES, and train on tumor database BRATS, then evaluate on Tumor2, as generalization baselines. Centralised training across all 5 databases (MultiUNet) is also shown. FedUniBrain is then trained with FL across 5 client databases (ATLAS, BRATS, MSSEG, TBI, WMH) using varying normalization methods, then evaluated on ISLES and Tumor2. The studied FL paradigm a) can reach performance similar to centralised training, even Oracles; b) generalizes to unseen databases better than task-specific models; c) IN and NF that don’t learn client-specific parameters generalize best.
@@
Method Mod. Drop Norm. Needs Targ. Data ISLES Tumor2 Average
BN Yes 48.5 75.9 61.9
IN Yes 53.2 76.5 64.9
Oracle: Disease specific models trained on target only NF Yes 54.8 73.4 64.1
Stroke-specific model (Trained on ATLAS (T1)) IN No 5.3 -
Tumor-specific model (Trained on BRATS) IN No - 69.1 37.2
Centralised (MultiUNet [41]) IN No 50.7 58.1 54.4
Centralised (MultiUnet [41]) IN No 55.5 72.2 63.9
FedUniBrain (avg. BN params) BN No 53.8 0.0 26.9
FedUniBrain (avg. BN params) BN No 54.5 68.0 61.3
FedUniBrain IN No 48.3 59.1 53.7
FedUniBrain IN No 55.3 72.7 64.0
FedUniBrain NF No 54.6 31.1 42.9
FedUniBrain NF No 52.8 72.1 62.5
FedUniBrain (client spec. BN params) BN Yes 50.2 36.8 43.5
FedUniBrain (client spec. BN params) BN Yes 49.9 70.1 60.0
Modality Drop for Robustness to Missing Modalities
Our desiderata is to train a model that can segment pathologies even in data with modality combinations different from those in the original training databases. This requires robustness against missing modalities. Results in Tab. 5 show that training with modality drop is effective for this purpose. When training with modality drop, the Dice score decreases by only 3.9% when evaluating with missing modalities compared to all modalities, versus 21.3% decrease without modality drop. Therefore, the following experiments focus on models trained with modality drop.
Table 5: Results from training FedUniBrain jointly on ATLAS, BRATS, MSSEG, TBI, and WMH using client-specific BN layers, with and without modality drop. Evaluation is done on the same databases with all modalities and by randomly dropping modalities during testing (missing modalities). Training with modality drop enables good performance with different sets of modalities at test time. Average Dice score over all databases is shown.
Training Method Evaluation (Dice (%)) $\uparrow$ $\Delta$ Dice $\downarrow$
All Modalities Missing Modalities
No Modality Drop 68.7 47.4 21.3
Modality Drop 69.1 65.2 03.9
Figure 4: Impact of a new client joining the federation. Performance (Dice score %) for both the initially trained clients and a client that joins at a later stage of training is presented. Blue bars represent the results before the new client joins (if the new client has an unseen modality, it is excluded during evaluation for the blue bar results). (a) Model initially trained on ATLAS, BRATS, MSSEG, TBI, and WMH (blue), then two scenarios are presented: after integrating ISLES client and learning its data (orange), and after integrating Tumor2 client (green). (b) Initial model trained without TBI (blue) and after integration and learning the TBI client (orange). (c) Initially trained without MSSEG (blue) and after integrating MSSEG (orange). Single-center training results on the new client are shown for comparison (violet).
Zero-Shot Generalization to Unseen Modality Combinations
We here assess the capabilities of generalizing to unseen databases with modality combinations not seen during training by evaluating on ISLES and Tumor2. Related results are shown in Tab. 4. Remarkably, despite not having seen any data from ISLES or Tumor2 during training, the best federated model’s average Dice score is only 0.9% lower than the Oracle models, which were trained on the specific target databases. FedUniBrain trained with FL across 5 clients reaches performance levels of centralized training using the 5 databases (MultiUnet, with IN as it performed best), demonstrating feasibility of decentralised training with such diverse brain lesion MRI databases.
Traditionally, models are trained on databases specific to a disease and evaluated on unseen clients with the same disease, as we trained with ATLAS (for stroke lesion) and evaluated on ISLES (stroke database), and trained with BRATS (for brain tumors) and evaluated on Tumor2 (tumor database), setting this as baseline. Both are only evaluated using T1 since it is the only modality common between training and test databases. For the BRATS model, we trained with modality drop to allow inference on just T1 in the Tumor2 database. The remaining non-Oracle results for ISLES are evaluated on all ISLES modalities except DWI, as DWI is unique to ISLES and was not seen during training. Federated joint training achieves a performance gain of 47.5 on ISLES and 3.0 on Tumor2, reinforcing the benefits of joint training across multiple MRI databases compared to pathology specific models.
Training with modality drop proves effective in dissociating specific modality combinations with certain pathologies. FedUniBrain with modality drop tends to show improved average performance over the two new databases with modality combinations not seen during training. Visual results are shown in Fig. 3(b). There is a large improvement for Tumor2 in all settings. Similarly, for ISLES, there is a general trend of improvement in most cases. In the few instances where the federated models performed slightly worse with modality drop, it may be due to stochasticity in federated optimization.
Non-client specific feature normalization, particularly IN, gives the best generalization results, with NF giving the second-best results. Therefore, these approaches are preferable for generalization compared to averaging client-specific BN layers, and estimating running statistics from target data, which yield the worst results. Additionally, the estimation requires unlabeled data the from unseen client.
Effect of Feature Normalization on Generalization
@@
Results reported in Tab. 2 indicate that learning client-specific parameters of BN yields better performance when evaluating on data from source clients, compared to other normalization methods. As shown in Tab. 4, however, when evaluating generalization on unseen databases (ISLES, Tumor2), BN with parameters tailored to source-clients underperforms compared to IN and NF. Consequently, a trade-off must be made when designing a model for training with this type of heterogeneous data: either prioritize performance on source clients that contributed training data or generalize to unseen databases.
Client with new modality and pathology joins federation
We evaluate the flexibility of our FedUniBrain framework in incorporating a new client that joins the federation at a later stage of training to further enhance the model with knowledge from new data. We consider three scenarios. First, a client with a modality combination not previously seen during training joins. Initially, the model is trained on ATLAS, BRATS, MSSEG, TBI, and WMH, then Tumor2 database joins (Fig. 4(a)). Second, a client with a modality not seen during earlier training joins, specifically the ISLES database that includes the unseen DWI modality (Fig. 4(a)). Lastly, a client with a modality and pathology not seen during training joins. We test this scenario in two cases: excluding TBI from initial training, the only database with SWI modality and Traumatic Brain Injuries (Fig. 4(b)), and excluding MSSEG, which is the only database with PD modality and Multiple Sclerosis (Fig. 4(c)).
The FedUniBrain model is trained for 300 communication rounds before a new client joins, followed by 200 additional rounds (optimization as per 4.2). Results in Fig. 4 show that the Dice scores for the 5 initial training databases remain stable after a new client joins and even slightly improve when a client with a previously unseen pathology joins (Fig. 4(b)4(c)). Moreover, FedUniBrain learns to segment the new modalities and pathologies of new clients, outperforming training only on the specific client (single-center), which indicates positive knowledge transfer. Even in the challenging case when a client with a previously unseen modality joins, our framework handles it without training re-initialization or complex modifications, enabling seamless continuation of training. Experiments on zero-shot generalization (Tab. 4) showed strong performance on unseen databases with pathologies seen during training. However, as shown by the blue bars in Fig. 4(b) and 4(c), generalization struggles with pathologies not seen during training. The FedUniBrain framework can easily integrate new clients, even with previously unseen pathologies and modalities, and learn to effectively segment the new type of data.
@@ 5 Conclusion
This study demonstrated the feasibility and effectiveness of using Federated Learning to train a single segmentation model across decentralised brain MRI databases, each with different input modalities and pathologies. Our results show that federated joint training on multiple databases results in a single model that can segment all diseases and sets of modalities seen in training clients with performance similar to client-specific models tailored for a specific set of modalities and disease. Promising generalization to unseen clients with modality combinations not seen during training is also demonstrated. The study finally highlights the importance of choosing feature normalization method, revealing a possible trade-off between generalization to unseen databases or tailored performance for clients that contribute to training. Having demonstrated, for the first time, the feasibility of FL from clients with such diverse brain lesion MRI data, the study opens the possibility for real-world collaborative federations to be initiated for this application.